Compound amount


accumulate 300,000 for eight years at 7% compounded quarterly. How much is the interest?



1
Expert's answer
2021-03-18T06:00:53-0400
"A=P(1+\dfrac{r}{n})^{n(t)}"

"I=A-P"

Given

"A=300,000"

"r=0.07"

"n=4"

"t=8"


"P=\dfrac{A}{(1+\dfrac{r}{n})^{n(t)}}"

"I=A-\dfrac{A}{(1+\dfrac{r}{n})^{n(t)}}"

"I=300000-\dfrac{300000}{(1+\dfrac{0.07}{4})^{4(8)}}=127805.26"